# Optimizing the interaction geometry of inverse Compton scattering x-ray sources

### v1

@@ II.2 Focused laser pulse
Before the diffraction limit is reached, the interaction time between the electron and laser pulse can become significantly reduced. This occurs when the Rayleigh range becomes similar to the laser pulse length. Not only does this impact the the total number of scattered x-rays, the energy spread is also increased. This effect is illustrated in Fig. 3. To analyze the effect of focused laser pulse we now consider the photon density of a $\mathrm{TEM}_{00}$ pulsed Gaussian laser beam
$\rho_{L}=\frac{N_{L}}{(2\pi)^{3/2}\sigma^{2}_{L}(z)\ \sigma_{t}}\exp\left[-\frac{x^{2}+y^{2}}{2\sigma_{L}^{2}(z)}-\frac{(z/c+t)^{2}}{2\sigma_{t}^{2}}\right].$ (11)
Figure 3: Illustration of how the laser divergence effects the interaction time when the pulse length becomes comparable to the interaction length. The bottom graph shows the product of the electron and laser density as a function of time for a tight and loose focus.
Where $\sigma_{t}$ is the laser pulse length and $\sigma_{L}(z)=\sigma_{L}(0)\sqrt{1+(z/z_{R})^{2}}$ with $z_{R}=2k_{L}\sigma_{L}^{2}$ the Rayleigh range. Throughout this paper we assume perfect laser beams ($M^{2}=1$). For non-perfect laser beams the analysis still holds, however, the Rayleigh length and angular spread should be adjusted accordingly. For now, we assume the electron beam to be described by a pencil beam. Integration of Eq. (6) with (11) results in
@@
Here, $\mathrm{erfcx}(x)=\exp(x^{2})\mathrm{erfc}(x)$ is the scaled complementary error function [abramowitz1948handbook], and $u=2z_{R}/c\sigma_{t}$. The term $\sqrt{\pi}\ u\ \mathrm{erfcx}(u)$ describes how the effective interaction time is reduced due to the laser divergence [hartemann2013nonlinear]. If we assume $\epsilon_{n}/\gamma\ll\lambda_{L}/4\pi$, a very reasonable assumption in practice, then also the divergence and finite longitudinal size of the electron beam, $\sigma_{z}$, can be taken into account by using $u=2z_{R}/\sqrt{(c^{2}\sigma_{t}^{2}+\sigma_{z}^{2})}$.
@@
To calculate the brilliance, we must now take into account the x-ray bandwidth, to which both the laser pulse and the electron beam contribute. Here we consider only the contributions of the laser pulse which has three sources. First, the Fourier limited bandwidth $\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\omega_{L}}=1/(2\omega_{L}\sigma_{t})$. However, due to the shorter interaction time, this term is broadened as
$\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\omega_{L}}=1/\left[2\omega_{L}\sigma_{t}\sqrt{\pi}\ u\ \mathrm{erfcx}(u)\right].$ (13)
@@
Second, the angular spread of a focused laser pulse gives rise to a spread in the interaction angle $\sigma_{\theta L}=1/(2k_{L}\sigma_{L})$, resulting in an additional x-ray broadening. By expanding Eq. (1) around $\theta_{L}=\pi$ we find $\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta L}}=1/\left(2k_{L}\sigma_{L}\right)^{2}$ [petrillo2012photon], which can be expressed in terms of $u$ as
$\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta L}}=1/(\omega_{L}\sigma_{t}u)$ (14)
@@
Finally, at high field intensities the x-ray energy spread is increased due to nonlinear broadening [hartemann2010low], characterized by the normalized vector potential amplitude $A_{0}=eE_{0}/(mc\omega_{L})$, with $e$ the elementary charge and $m$ the electron mass. The corresponding x-ray energy spread is [petrillo2012photon]
@@
$\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{A_{0}}\simeq\frac{1}{6}A_{0}^{2}=\chi/(\omega_{L}\sigma_{t}u).$ (15)
@@
In the right equality, we have expressed the nonlinear broadening in terms of the laser pulse dimensions. Here, $\chi=4\sigma_{T}N_{L}/(3\sqrt{\pi}\alpha\lambda_{L}c\sigma_{t})$ relates the laser peak power to the normalized vector potential and $\alpha$ is the fine structure constant. The nonlinear broadening can become the significant contribution to the x-ray energy spread at tight laser focus. We note that the nonlinear broadening can be circumvented using appropriate laser pulse chirping [terzic2021laser].
@@
The three energy spread contributions can be combined, resulting in the following expression for x-ray brilliance
$B_{\mathrm{x}}=\frac{3f\sigma_{\mathrm{T}}N_{L}N_{\mathrm{e}}}{4\pi\sigma_{e}^{2}}\frac{\gamma^{2}\omega_{L}\sigma_{t}}{\pi^{2}\sigma_{e}^{2}}\frac{\pi u^{2}\ \mathrm{erfcx}(u)^{2}}{\sqrt{1+4\pi(1+\chi^{2})\mathrm{erfcx}(u)^{2}}}.$ (16)
@@
To optimize the brilliance, we note that for space charge limited electron beams $\epsilon_{n}=\eta\sqrt{eN_{e}}$ [pasmans2016extreme], where $\eta$ is a property of the injector that describes the electron beam emittance at given bunch charge. Therefore $N_{e}/\sigma_{e}^{2}=\mathrm{constant.}$ Since $u\propto\sigma_{L}^{2}=\sigma_{e}^{2}$, the function $g(u)=\pi u\ \mathrm{erfcx}(u)^{2}/\sqrt{1+4\pi(1+\chi^{2})\mathrm{erfcx}(u)^{2}}$ needs to be maximized, resulting in the optimized value for $\sigma_{e}$ that maximizes the x-ray brilliance
@@
where $u_{\mathrm{max}}$ is the value of $u$ that maximizes $g(u)$. To take into account the electron pulse length $\sigma_{z}$, the substitution $\sigma_{t}\rightarrow\sigma_{t}\sqrt{1+(\sigma_{z}/c\sigma_{t})^{2}}$ should be used in Eq. (17). To take into account the bandwidth contributions from the electron beam, additional terms should be added to the square root of $g(u)$ which lowers the value in $u_{\mathrm{max}}$.
The green and blue curves in Fig. 4 show the functions $g(u)$ for $\chi=0$ and $\chi=5$, respectively. The inset show $u_{\mathrm{max}}$ as a function of $\chi$ The yellow curve shows the function $1/u$, which corresponds to the brilliance scaling of the pencil beam limit.
This analysis has been done under the assumption that the electron and laser spot sizes are matched, $\sigma_{e}=\sigma_{L}$. At a fixed energy, the electron spot size is determined by the beam divergence $\sigma_{\theta e}$ and normalized emittance $\epsilon_{n}$. Due to practical considerations, for relativistic beams, $\sigma_{\theta e}$, is typically limited a few milliradians. We therefore consider a fixed electron divergence and optimize the normalized emittance such that $\sigma_{e}=\sigma_{L}$. To arrive at an expression for the optimized bunch charge we use the fact that $\epsilon_{n}=\eta\sqrt{eN_{e}}=\gamma\sigma_{e}\sigma_{\theta e}$, combined with Eq. (17). This results in
@@
Figure 4: Plots of the function $g(u)$ for $\chi=0$ (green) and $\chi=5$ (blue). The inset shows value of $u_{\mathrm{max}}$ for varying values of $\chi$. The green curve shows the function $1/u$ corresponding to the brilliance scaling of the pencil beam limit.
@@ II.3 Optimized head-on geometry
As an example, we calculate the brilliance of an ICS source generating 10 keV x-rays in a head-on geometry. The electron bunches are generated from a C-band photoinjector [lucas2023toward] that is able to create 200 pC electron bunches at $\epsilon_{n}=200\mathrm{\ nmrad}\ (\eta=14\mathrm{\ nmrad/\sqrt{pC}})$ and a relative energy spread of 0.2%. The electrons are accelerated to $\gamma=46$ to generate 10 keV x-rays with a 5 mJ, 100 fs, 1030 nm interaction laser pulse [LightConversion_Pharos]. We assume the electron pulse length to be matched to the laser pulse length. Both the photoinjector and laser system operate at $f=1$ kHz.
Figure 5 shows the x-ray brilliance at a varying bunch charge where it is assumed that $\sigma_{L}=\sigma_{e}=\epsilon_{n}/(\gamma\sigma_{\theta})$. Here we have also included the energy spread contributions from the electron beam. The continuous curve is calculated using Eq. (16), and the dots are the result of particle tracking simulations (General Particle Tracer [PulsarPhysicsGeneralGPT]) and subsequent numerical calculation of the Liénard-Wiechert potentials [thomas2010algorithm]. At these parameters $\chi=5.6$ and $\sigma_{e}^{\mathrm{opt}}=2.4\ \mathrm{\mu m}$. With an electron divergence angle of $\sigma_{\theta}=1$ mrad, the optimized bunch charge is 43 pC.
At the relatively low bunch charge of 43 pC, the x-ray brilliance is 10 times higher than at 1 nC. Additionally, the total number of scattered x-rays is only 4% lower at 43 pC than at 1 nC. Due to the optimization of the laser focus and bunch charge, ICS becomes significantly more efficient at 20 times lower bunch charge. This can be a very important design consideration regarding background suppression and radiation protection.
Figure 5: X-ray brilliance in for 10 keV x-ray generation in a head-on geometry for varying bunch charge and matched electron and laser spot sizes. The analytical result of Eq. (16) (solid curve) agrees well with simulation results (dots). The arrow marks the theoretical optimized bunch charge according to Eq. (18)
@@
In this section we investigate how to make use of the fact that a Compton x-ray source becomes more efficient at high electron energies. In a head-on geometry, the desired x-ray energy determines electron energy. However, if instead the interaction angle is varied to achieve the required x-ray energy, the favorable scaling with electron beam energy can be exploited even further. We will find that this geometry can be particularly useful for soft x-ray generation
We will now use the framework of covariant electrodynamics, with the metric, $g^{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$, to describe the interaction of a relativistic electron bunch with a laser pulse under an arbitrary interaction angle. We start by considering a single electron interacting with a Gaussian pencil beam. Next, we add the effects of a focused laser beam and describe the electron bunch as a finite sized object. In Sec.III.4 the case where $\theta_{L}\ll 1$ is considered and the x-ray brilliance is optimized for this grazing angle geometry. Finally, a comparison is made between grazing angle and head-on ICS.
@@ III.1 Single electron with a pencil laser beam
In the regime where electron recoil can be ignored, i.e. $\hbar\omega_{\mathrm{x}}/\gamma m_{e}c^{2}\ll 1$, a single electron with 4-trajectory $x^{\mu}(\tau)$ and 4-velocity $u^{\mu}(\tau)$ radiates, according to the Liénard-Wiechert potentials, with a spectral angular density [hartemann2013nonlinear]
@@
Here $k_{\mathrm{x}}^{\mu}=k_{\mathrm{x}}(1,\bm{n}_{\mathrm{x}})$ is the four- wave vector of the scattered x-ray with $\bm{n}_{\mathrm{x}}$ its propagation direction. $x^{\mu}=(ct,\bm{x})$ is the electron four-position, $u^{\mu}=\gamma(1,\bm{\beta})$ its normalized four-velocity, and $\tau=t/\gamma$ is the proper time.
The x-ray brilliance is closely related to the spectral angular density. In the approximation $\sigma_{e}\ll\sigma_{L}$,
$B_{\mathrm{x}}\simeq\frac{fN_{e}}{\pi\sigma_{e}^{2}}\omega_{\mathrm{x}}\mathcal{S}(\omega_{\mathrm{x}},0).$ (20)
The dimensionless quantity $\omega_{\mathrm{x}}\mathcal{S}(\omega_{\mathrm{x}},0)$ is the peak normalized spectral angular density, which is maximized at $\theta_{\mathrm{x}}=0$ and when $\omega=\omega_{\mathrm{x}}$. In this section we will analyze how $\omega_{\mathrm{x}}\mathcal{S}(\omega_{\mathrm{x}},0)$ depends on the interaction geometry of ICS. Specifically, how it depends on the laser pulse dimensions, interaction angle and electron energy.
The integral in Eq. (19) is solved by considering a single electron that experiences the normalized vector potential, with amplitude $A(\tau)$, of the laser pulse. We assume that this normalized vector potential is Gaussian, i.e. $A(\tau)=A_{0}\exp(-\tau^{2}/2\tau_{\mathrm{int}}^{2})$. Here, $\tau_{\mathrm{int}}$ is the proper interaction time of the electron with the laser pulse. Note that $\tau_{\mathrm{int}}$ does not equal the laser pulse length. How $\tau_{\mathrm{int}}$ relates to the laser pulse parameters is discussed later. $A_{0}$ can be written in terms of the laser pulse dimensions
@@
Due to the asymmetry in the interaction geometry, we also consider an asymmetric laser pulse focus with two transverse sizes, $\sigma_{L\parallel}$ and $\sigma_{L\perp}$. Here, the subscripts $\parallel$ and $\perp$ refer to the directions parallel and perpendicular to the x-z plane, respectively.
When nonlinear effects can be neglected, i.e. $A_{0}\ll 1$, the peak normalized spectral angular density from a single electron is given by the concise expression
$\omega_{\mathrm{x}}\mathcal{S}(\omega_{\mathrm{x}},0)$ $=\frac{1}{8\pi}\alpha A_{0}^{2}\omega_{\mathrm{x}}^{2}\tau_{\mathrm{int}}^{2}$ (22)
For details, we refer to Appendix A. In this expression all dependencies on the interaction geometry are implicit in $A_{0}$ and $\tau_{\mathrm{int}}$. In a head-on geometry ($\theta_{L}=\pi$) the proper interaction time $\tau_{\mathrm{int}}=c\sigma_{t}/[\gamma(1+\beta)]$, while for $\theta_{L}=0$, it becomes $\tau_{\mathrm{int}}=c\sigma_{t}/[\gamma(1-\beta)]$.
@@
For intermediate interaction angles, $0<\theta_{L}<\pi$, $\tau_{\mathrm{int}}$ depends on $\sigma_{L\parallel}$. This effect is explained in Fig. 6a, in which the electron (green) crosses the laser pulse (red) under an angle $\theta_{L}$. The vector potential amplitude, as experienced by the electron, is illustrated by the red solid curve. Due to the finite value of $\sigma_{L\parallel}$, the time that electron and laser pulse overlap is reduced. In the co-moving frame (Fig. 6b) this can be understood as the electron experiencing a tilted laser pulse front. Therefore, the experienced vector potential amplitude becomes a function of both the pulse length, $\sigma_{t}$ and transverse size $\sigma_{L\parallel}$.
The reduction in the interaction time can be quantified by a dimensionless parameter $\zeta\in[0,1]$, defined as the ratio between the proper interaction time and the Lorentz transformed laser pulse length, and is given by
$\zeta=\left[1+\left(\frac{\beta\sin\theta_{\mathrm{L}}}{1-\beta\cos{\theta_{L}}}\frac{c\sigma_{t}}{\sigma_{L\parallel}}\right)^{2}\right]^{-1/2}.$ (23)

### v2

@@ II.2 Focused laser pulse
Before the laser beam diffraction limit is reached, the Rayleigh range may become comparable to, or even smaller than, the laser pulse length, resulting in a significant reduction of the interaction time between the electron pulse and the laser pulse (see Fig. 3). This will lead to both a reduction of the number of x-ray photons produced, and an increase of the x-ray energy spread. To analyze the effect of focusing the laser beam we now consider the photon density of a $\mathrm{TEM}_{00}$ pulsed Gaussian laser beam
$\rho_{L}=\frac{N_{L}}{(2\pi)^{3/2}\sigma^{2}_{L}(z)\ \sigma_{t}}\exp\left[-\frac{x^{2}+y^{2}}{2\sigma_{L}^{2}(z)}-\frac{(z/c+t)^{2}}{2\sigma_{t}^{2}}\right],$ (11)
Figure 3: Illustration showing how focusing the laser beam results in a reduction of the Rayleigh length, which affects the interaction time. The bottom graph shows the product of the electron pulse density and laser pulse photon density as a function of time for a tight focus and a loose focus.
where $\sigma_{t}$ is the laser pulse length and $\sigma_{L}(z)=\sigma_{L}(0)\sqrt{1+(z/z_{R})^{2}}$ with $z_{R}=2k_{L}\sigma_{L}^{2}$ the Rayleigh length. Throughout this paper we assume perfect laser beams ($M^{2}=1$). For non-perfect laser beams the analysis still holds if the expressions for the Rayleigh length and the angular spread are adjusted accordingly. For now, we assume the electron beam to be described by a pencil beam. Substituting (11) into the central term of Eq. (7) and subsequently performing the integration over time results in
@@
Here, $\mathrm{erfcx}(x)=\exp(x^{2})\mathrm{erfc}(x)$ is the scaled complementary error function [1], and $u=2z_{R}/c\sigma_{t}$. The term $\sqrt{\pi}\ u\ \mathrm{erfcx}(u)$ describes how the effective interaction time is reduced due to laser beam divergence [21]. If we assume $\epsilon_{n}/\gamma\ll\lambda_{L}/4\pi$, a very reasonable assumption in practice, then also the divergence and finite longitudinal size $\sigma_{z}$ of the electron beam can be taken into account using $u=2z_{R}/\sqrt{(c^{2}\sigma_{t}^{2}+\sigma_{z}^{2})}$.
@@
To calculate the brilliance, we must now take into account the x-ray energy spread, to which both the laser beam properties and the electron beam properties contribute. Here we consider only the three contributions of the laser beam. First, we take into account the Fourier limited bandwidth $\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\omega_{L}}=1/(2\omega_{L}\sigma_{t})$, which is broadened due to the shorter interaction time:
$\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\omega_{L}}=1/\left[2\omega_{L}\sigma_{t}\sqrt{\pi}\ u\ \mathrm{erfcx}(u)\right].$ (13)
@@
Second, the angular spread of a focused laser pulse gives rise to a spread in the interaction angle $\sigma_{\theta L}=1/(2k_{L}\sigma_{L})$, resulting in additional x-ray broadening. By expanding Eq. (2) around $\theta_{L}=\pi$ we find $\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta L}}=1/\left(2k_{L}\sigma_{L}\right)^{2}$ [29], which can be expressed in terms of $u$ as
$\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta L}}={1\over\omega_{L}\sigma_{t}u}$ (14)
@@
Finally, at high field intensities the x-ray energy spread is increased due to nonlinear broadening [22], characterized by the normalized vector potential amplitude $A_{0}=eE_{0}/(mc\omega_{L})$. The corresponding x-ray energy spread is given by [29]
@@
$\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{A_{0}}\simeq\frac{1}{6}A_{0}^{2}={\chi\over\omega_{L}\sigma_{t}u}.$ (15)
@@
In the right hand side term we have expressed the nonlinear broadening in terms of the laser pulse dimensions. Here, $\chi=4\sigma_{T}N_{L}/(3\sqrt{\pi}\alpha\lambda_{L}c\sigma_{t})$ relates the laser peak power to the normalized vector potential and $\alpha$ is the fine structure constant. $\chi$ is only dependent on the laser pulse energy and duration, and not the focal spot size. The nonlinear broadening can become the significant contribution to the x-ray energy spread if the laser is tightly focused. We note that the nonlinear broadening can be circumvented using appropriate laser pulse chirping [40], which would effectively bring $\chi\rightarrow 0$.
@@
The three energy spread contributions associated with the laser beam properties can be combined, resulting in the following expression for the x-ray brilliance
$B_{\mathrm{x}}=\frac{3f\sigma_{\mathrm{T}}N_{L}N_{\mathrm{e}}}{\sqrt{2\pi}\ 8\pi^{3}}\frac{\gamma^{2}\omega_{L}\sigma_{t}}{\sigma_{e}^{4}}\frac{\pi u^{2}\ \mathrm{erfcx}(u)^{2}}{\sqrt{1+4\pi(1+\chi^{2})\mathrm{erfcx}(u)^{2}}}.$ (16)
@@
To optimize brilliance, we note that for space charge limited electron beams $\epsilon_{n}=\eta\sqrt{eN_{e}}$ [28], where $\eta$ is a property of the injector that describes the electron beam emittance at given bunch charge. Therefore $N_{e}/\sigma_{e}^{2}=\mathrm{constant.}$ Since $u\propto\sigma_{L}^{2}=\sigma_{e}^{2}$, the function $g(u)=\pi u\ \mathrm{erfcx}(u)^{2}/\sqrt{1+4\pi(1+\chi^{2})\mathrm{erfcx}(u)^{2}}$ must be maximized, resulting in the optimized value for $\sigma_{e}$
@@
where $u_{\mathrm{max}}$ is the value of $u$ that maximizes $g(u)$ and thus the x-ray brilliance. To take into account the length of the electron pulse $\sigma_{z}$, the substitution $\sigma_{t}\rightarrow\sigma_{t}\sqrt{1+(\sigma_{z}/c\sigma_{t})^{2}}$ should be used in Eq. (17). If bandwidth contributions from the electron beam are taken into account, additional terms should be added to the square root of $g(u)$. Then, a larger energy spread is allowed from the laser pulse, resulting in a lower value of $u_{\mathrm{max}}$.
In Fig. 4 the function $g(u)$ is plotted for $\chi=0$ (green curve) and $\chi=5$ (blue curve), respectively. The function $1/u$ (yellow curve) is plotted as well, which gives the brilliance scaling in the pencil beam limit. The inset shows a plot of $u_{\mathrm{max}}$ as a function of $\chi$.
This analysis has been performed under the assumption that the electron and laser spot sizes are matched, i.e., $\sigma_{e}=\sigma_{L}$. For a fixed energy, the electron spot size is determined by the beam divergence $\sigma_{\theta e}$ and the normalized emittance $\epsilon_{n}$. For relativistic beams $\sigma_{\theta e}$ is typically limited in practice to a few milliradians. We choose a fixed electron beam divergence of 1 mrad and optimize the normalized emittance such that $\sigma_{e}=\sigma_{L}$. To arrive at an expression for the optimized bunch charge we use the fact that $\epsilon_{n}=\eta\sqrt{eN_{e}}=\gamma\sigma_{e}\sigma_{\theta e}$, combined with Eq. (17). This results in
@@
Figure 4: Plots of the function $g(u)$ for $\chi=0$ (green curve) and $\chi=5$ (blue curve). The inset shows a plot of $u_{\mathrm{max}}$ as a function of $\chi$. For comparison the function $1/u$ (yellow curve) is plotted as well, which gives the brilliance scaling of the pencil beam limit.
@@ II.3 Optimized head-on geometry
As an example, we calculate the brilliance of an ICS source generating 10 keV x-rays in a head-on geometry. The electron bunches are generated from a C-band photoinjector [26, 2] that is able to create 200 pC electron bunches at $\epsilon_{n}=200\mathrm{\ nmrad}\ (\eta=14\mathrm{\ nmrad/\sqrt{pC}})$ and a relative energy spread of 0.2%. The electrons are accelerated to $\gamma=46$ to generate 10 keV x-rays with a 5 mJ, 100 fs, 1030 nm interaction laser pulse ($\chi=5.6$) [30]. In this example, we take the electron divergence angle to be $\sigma_{\theta}=1$ mrad, and the electron pulse length and the laser pulse length are equal. Both the photoinjector and the laser operate at $f=1$ kHz.
Figure 5 shows the x-ray brilliance as a function of bunch charge. Here the bunch charge determines $\sigma_{e}=\eta\sqrt{eN_{e}}/(\gamma\sigma_{\theta e})$ and the size of the laser is matched to the electron beam, $\sigma_{L}=\sigma_{e}$. This results in a bunch charge where $\sigma_{L}$ and $\sigma_{e}$ are optimized. The energy spread contributions due to the electron beam properties have also been included. The continuous curve is calculated using Eq. (16), and the dots are the result of particle tracking simulations (General Particle Tracer [32]) and subsequent numerical calculation of the Liénard-Wiechert potentials [41]. From Eqs. 17 and 18 it then follows that $\sigma_{e}^{\mathrm{opt}}=2.4\ \mathrm{\mu m}$ and the corresponding optimized bunch charge $eN_{e}^{\mathrm{opt}}=43\,{\rm pC}$.
At the relatively low bunch charge of 43 pC, the x-ray brilliance is 10 times higher than at 1 nC. Additionally, the total number of scattered x-rays is only 4% lower at 43 pC than at 1 nC. Remarkably, due to the optimization of the laser focus and bunch charge, ICS is significantly more efficient at a bunch charge which is $\sim$20 times lower.
Figure 5: X-ray brilliance for 10 keV x-ray generation in a head-on geometry with matched electron and laser spot sizes, as a function of bunch charge. The analytical result of Eq. (16) (solid curve) agrees well with simulation results (dots). The arrow marks the optimized bunch charge given by Eq. (18)
@@
In a head-on ICS geometry, the generated x-ray energy increases with the electron energy. For a given laser wavelength, the desired x-ray energy therefore determines the electron energy. At the same time, the x-ray brilliance increases with electron energy, implying that, using ICS, the generation of hard x-rays is more favorable than the generation of soft x-rays. However, if instead of the electron energy, the interaction angle is varied to achieve the required x-ray energy, an additional degree of freedom is introduced, which allows further exploitation of the favorable scaling of the brilliance with the electron beam energy. This will be explored in some detail in this section. We will find that a grazing angle geometry (see Fig. 1) can be particularly useful for the generation of high-brilliance soft x-rays.
We will now use the framework of covariant electrodynamics, with the metric $g^{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$, to describe the interaction of a relativistic electron bunch with a laser pulse under an arbitrary interaction angle. We start, in Sec. III.1, by considering a single electron interacting with a Gaussian pencil beam. Next, in Sec. III.2, we investigate the effects of focusing the laser beam, and in Sec. III.3, an electron bunch of finite dimensions instead of a single electron. In Sec. III.4 the x-ray brilliance is optimized for $\theta_{L}\ll 1$. Finally, in Sec. III.5, the optimized performances of ICS in a grazing angle geometry and in a head-on geometry are compared.
@@ III.1 Single electron with a pencil laser beam
In the regime where electron recoil can be ignored, i.e., $\hbar\omega_{\mathrm{x}}/\gamma mc^{2}\ll 1$, a single electron with 4-trajectory $x^{\mu}(\tau)$ and 4-velocity $u^{\mu}(\tau)$ radiates, according to the Liénard-Wiechert potentials, with a spectral angular density [21]
@@
Here $k_{\mathrm{x}}^{\mu}=k_{\mathrm{x}}(1,\bm{n}_{\mathrm{x}})$ is the four-wave vector of the scattered x-rays with $\bm{n}_{\mathrm{x}}$ the propagation direction. Furthermore, $x^{\mu}=(ct,\bm{x})$ is the electron four-position, $u^{\mu}=\gamma(1,\bm{\beta})$ its normalized four-velocity, and $\tau=t/\gamma$ is the proper time.
The x-ray brilliance is closely related to the spectral angular density. Using the fact that, obviously, $\sigma_{e}\ll\sigma_{L}$ for a single electron,
$B_{\mathrm{x}}\simeq\frac{fN_{e}}{2\pi\sigma_{e}^{2}}\omega_{0}\mathcal{S}(\omega_{0},0).$ (20)
The dimensionless quantity $\omega_{0}\mathcal{S}(\omega_{0},0)$ is the peak normalized spectral angular density, which is maximized for $\theta_{\mathrm{x}}=0$ and therefore $\omega=\omega_{0}=2\gamma^{2}\omega_{L}(1-\beta\cos{\theta_{L}})$. In this section we will analyze how $\omega_{0}\mathcal{S}(\omega_{0},0)$ depends on the interaction geometry of ICS. Specifically, how it depends on the laser pulse dimensions, interaction angle and electron energy.
The integral in Eq. (19) is solved by considering a single electron that experiences the normalized vector potential, with amplitude $A(\tau)$, due the laser pulse. We assume that this normalized vector potential is Gaussian, i.e. $A(\tau)=A_{0}\exp(-\tau^{2}/2\tau_{\mathrm{int}}^{2})$. Here, $\tau_{\mathrm{int}}$ is the proper interaction time of the electron with the laser pulse. Note that $\tau_{\mathrm{int}}$ is not equal to the length of the laser pulse. The manner in which $\tau_{\mathrm{int}}$ depends on the parameters of the laser pulse is discussed later. $A_{0}$ can be written in terms of the laser pulse dimensions
@@
Because of the asymmetry in the interaction geometry, we also have to consider an asymmetric laser pulse focus with two transverse sizes, $\sigma_{L\parallel}$ and $\sigma_{L\perp}$. Here, the subscripts $\parallel$ and $\perp$ refer to the directions parallel and perpendicular to the x-z plane, respectively.
When nonlinear effects can be neglected, i.e. $A_{0}\ll 1$, the peak normalized spectral angular density due to a single electron is given by the concise expression
$\omega_{\mathrm{0}}\mathcal{S}(\omega_{\mathrm{0}},0)$ $=\frac{1}{8\pi}\alpha A_{0}^{2}\omega_{\mathrm{x}}^{2}\tau_{\mathrm{int}}^{2}$ (22)
For details, we refer to Appendix A. In this expression all dependencies on the interaction geometry are implicit in $A_{0}$ and $\tau_{\mathrm{int}}$. In a head-on geometry ($\theta_{L}=\pi$) the proper interaction time $\tau_{\mathrm{int}}=\sqrt{2}\sigma_{t}/[\gamma(1+\beta)]$, while for $\theta_{L}=0$, $\tau_{\mathrm{int}}=\sqrt{2}\sigma_{t}/[\gamma(1-\beta)]$.
@@
For intermediate interaction angles, $0<\theta_{L}<\pi$, $\tau_{\mathrm{int}}$ depends on $\sigma_{L\parallel}$. This is illustrated in Fig. 6a, which shows the electron (green) crossing the laser pulse (red) under an angle $\theta_{L}$. The vector potential amplitude, as experienced by the electron, is illustrated by the red solid curve. Due to the less than infinite value of $\sigma_{L\parallel}$, the time that the electron and the laser pulse overlap is reduced; for $\sigma_{L\parallel}=0$, the overlap time would be reduced to zero. In the co-moving frame (Fig. 6b) this can be understood as the electron experiencing a tilted laser pulse front. Therefore, the experienced vector potential amplitude becomes a function of both the pulse length, $\sigma_{t}$ and transverse size $\sigma_{L\parallel}$.
The reduction in the interaction time can be quantified by a dimensionless parameter $\zeta\in[0,1]$, defined as the ratio between the proper interaction time and the Lorentz transformed laser pulse length:
$\zeta\equiv\left[1+\left(\frac{\beta\sin\theta_{\mathrm{L}}}{1-\beta\cos{\theta_{L}}}\frac{c\sigma_{t}}{\sigma_{L\parallel}}\right)^{2}\right]^{-1/2}.$ (23)